Ceph/CVMFS/Filer Service Meeting

Europe/Zurich
600/R-001 (CERN)

600/R-001

CERN

15
Show room on map
    • 14:00 14:05
      CVMFS 5m
      Speaker: Enrico Bocchi (CERN)
      • bbp.epfl, na{49,61,62}.cern.ch migrated to a more powerful VM (still SLC6)
      • projects.cern.ch migrated to CC7
      • cvmfs-server-2.6.0 installed on sft-nightlies.cern.ch

       

      • cvmfs-server-2.6.1 has a bugfix for GC on na61.cern.ch. To be installed on the RM
      • RQF for new repo atlas-online-nightlies: CC7+S3 (no gateway)

       

      • lemon to collectd
      • review of RM configuration as per Jakob's comments
      • upgrade campaign to 2.6.0
      • review puppt manifest + split cvmfssrv
    • 14:05 14:10
      Ceph Upstream News 5m

      Releases, Tickets, Testing, Board, ...

      Speaker: Dan van der Ster (CERN)
    • 14:10 14:15
      Ceph Backends & Block Storage 5m

      Cluster upgrades, capacity changes, rebalancing, ...
      News from OpenStack block storage.

      Speaker: Theofilos Mouratidis (National and Kapodistrian University of Athens (GR))

      (Dan)

      • Ceph gabe: moved all data from default root to incoming root. (This moves all data off of filestore and onto Bluestore, and changes the crush tunables at the same time).
        • It has gone much more smoothly than expected!
        • Progressing ~20% per day -- 15% left.
        • Tuning:
          • ceph tell osd.* injectargs -- --osd_max_backfills=4 --osd_delete_sleep=2 --osd_target_transaction_size=15
      • Beesly:
        • Cluster is still quite imbalanced, but balancer doesn't find more changes. 12.2.12 might help, but more importantly we need to continue PG splitting. Currently at 5300 pgs -- heading slowly to 8192.
      • CEPH-701 ceph/gabe: zap then recreate old filestore osds
      • CEPH-702 Add RJ35/RJ37 to ceph/flax
        ​​
    • 14:15 14:20
      Ceph Disk Management 5m

      OSD Replacements, Liaison with CF, Failure Predictions

      Speaker: Julien Collet (CERN)

      Julien:

      • Improved disk replacement scripts to include cephwigner
    • 14:20 14:25
      S3 5m

      Ops, Use-cases (backup, DB), ...

      Speakers: Julien Collet (CERN), Roberto Valverde Cameselle (Universidad de Oviedo (ES))

      (Dan)

      • RQF1262152: the user is unable to test our patched rgw (12.2.11 0.7). I propose we upgrade all the rgw's and let him test.
        • If any problems, we roll back to v12.2.11 0.6.

      Julien

      • CEPH-697: everything ready to contact the users, will proceed today
    • 14:25 14:30
      CephFS/Manila/FILER 5m

      Filer Migration, CephFS/Manila status and plans.

      Speaker: Dan van der Ster (CERN)
    • 14:30 14:35
      HPC 5m

      Performance testing, HPC storage status and plans

      Speakers: Alberto Chiusole (Universita e INFN Trieste (IT)), Pablo Llopis Sanmillan (CERN)
    • 14:35 14:40
      HyperConverged 5m
      Speakers: Jose Castro Leon (CERN), Julien Collet (CERN), Roberto Valverde Cameselle (Universidad de Oviedo (ES))
    • 14:40 14:45
      Monitoring 5m
    • 14:45 14:50
      AOB 5m
      • Alberto last day Weds.
      • Dan absent Thursday until 29 Apr.

      Julien:

      • Meeting with UNIGE on Friday (Similar to the one with French Ministere de l'Interieur)
        • On the move to switch to Ceph, asked for advices
Your browser is out of date!

If you are using Internet Explorer, please use Firefox, Chrome or Edge instead.

Otherwise, please update your browser to the latest version to use Indico without problems.

×